OpenStudy (anonymous):

3^(1-2x)=4x need to put in log form and i am stuck thanks

OpenStudy (anonymous):

\[1-2x=\log_3(4x)\] but that will not help you to solve for x

OpenStudy (anonymous):

are you trying to get a number for x?

OpenStudy (anonymous):

no got to change from expoental to log

OpenStudy (anonymous):

ok then what i wrote is correct. the base on the left hand side it 3, to if you "take the log base 3" of both sides you get it

OpenStudy (anonymous):

ooooooooooh was it \[3^{1-2x}=4^x\]??

OpenStudy (anonymous):

thank you